OpenStudy (anonymous):

In the ellipse equation below, what are the coordinates of the center of the ellipse? (x + 10)2 (y - 8)2 ------- + ------ =1 49 25

jimthompson5910 (jim_thompson5910):

Hint: The center of the general ellipse below \[\Large \frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2} = 1\] is the point (h,k)